Show how two phase regions only need one variable to complete define the phase compositions as opposed to single phase fields which need two variables using the Gibb’s phase rule
Answer #1

Please find the attached solution.

Solution As Per Gibbs bhose rule P+F=C+2 where P= no. of phose F = Degree of freedom & no. of independent Variables to define